Rose bought green mangoes at P50.00 a kilo and sold them at 4 for P35.00. If she sold 20 kilos and there is an average of 7 mangoes per kilo, how much did she earn?
Answer #1
: Rose bought green mangoes at P50.00 a kilo and sold them at 4 for P35.00. If she sold 20 kilos and there is an average of 7 mangoes per kilo, how much did she earn? : Find the cost she paid for each Mango: {{{50/7}}} Find the no. of mangoes sold: 20 * 7 = 140 mangoes Find the price she sold each Mango: {{{35/4}}} : Let p = profit she made by buying and selling these Mangos : p = 140({{{(35/4) - (50/7)}}}) p = 140({{{(245/28) - (200/28)}}}) p = 140({{{45/28}}}) p = 5(45) p = 225 P profit on 20 kilos of Mangos
answered by: mokoena terry